OpenStudy (vshiroky):

Equations with infinite and no solutions: Solve each equation: 1) 6x+3-6x=3

OpenStudy (vshiroky):

My daughter put infinite... is that correct?

ganeshie8 (ganeshie8):

If there are infinitely many solutions, can you list a few of them ?

OpenStudy (vshiroky):

I dont even know what to do lol

ganeshie8 (ganeshie8):

You're given an equation

OpenStudy (vshiroky):

well the 6x cancels out when you combine them and so does the 3

ganeshie8 (ganeshie8):

A "number" is called a solution to an equation if you get a "true" statement when you plugin the "number" into the equation

ganeshie8 (ganeshie8):

6x+3-6x=3

ganeshie8 (ganeshie8):

maybe plugin x = 1 and see if "1" is a solution

ganeshie8 (ganeshie8):

6*1+3-6*1 = 3 6 + 3 - 6 = 3 3 = 3 which is true, therefore 1 is a solution to the given equation

OpenStudy (vshiroky):

so you just have to keep trying numbers

ganeshie8 (ganeshie8):

Nope, I have just given you an example to explain what is a solution

ganeshie8 (ganeshie8):

6x+3-6x=3 cancelling 6x both sides gives you 3 = 3 which is true no matter what x is. therefore "every number x" is a solution to the given equaiton

OpenStudy (vshiroky):

so that means it has to be infinite?

ganeshie8 (ganeshie8):

Yep, there are infinitely many numbers

OpenStudy (vshiroky):

so for instance 0=-4p+4p if you plug in 1 for both it still equals 0 so infintie?

ganeshie8 (ganeshie8):

Exactly! If you plugin 1 or 2 or 3 or anything for p, you still get a true statement : 0 = 0 therefore all numbers are solutions to above equation

OpenStudy (vshiroky):

ok what about a-a=-5

OpenStudy (vshiroky):

its no

ganeshie8 (ganeshie8):

a-a = -5 0 = -5 which is FALSE no matter what "a" is. Therefore there are NO solutions to the given equation

ganeshie8 (ganeshie8):

In other words, you can never find a value for "a" that satisfies above equation

OpenStudy (vshiroky):

ok great so i can try to plug in 1 for everything and if it doesnt work then there are none

OpenStudy (vshiroky):

and if it does there is infinite

ganeshie8 (ganeshie8):

not so simple, you need to convince yourself that there are indeed infinitely many solutions

ganeshie8 (ganeshie8):

But yeah, what you said works..